OpenStudy (anonymous):

Select ALL true statements about the graph of: This line has only an x-intercept. This equation does not form a line when graphed. This line has no intercepts. This line goes through the origin. This line has only a y-intercept. This line has both an x and y-intercept. This equation forms a line when graphed.

OpenStudy (mrnood):

I think you have missed the equation from your question....

OpenStudy (anonymous):

lol thanks :-D Select ALL true statements about the graph of: -x+4y=0 This line has only an x-intercept. This equation does not form a line when graphed. This line has no intercepts. This line goes through the origin. This line has only a y-intercept. This line has both an x and y-intercept. This equation forms a line when graphed.

OpenStudy (mrnood):

rewrite the equation in the form y = ....

OpenStudy (anonymous):

I have no idea how to do that.... :/

OpenStudy (mrnood):

are you sure you have no idea? it is an equation so bith sides reamin equal if oyu do the same thing to both sides. Try adding x to each side and rewrite the equation in your reply

OpenStudy (anonymous):

4y=x?

OpenStudy (mrnood):

yes Now divide both sides by 4

OpenStudy (anonymous):

y=x/4

OpenStudy (mrnood):

Good - so you DID know how to do it. So do you know about the equation of a line?

OpenStudy (anonymous):

Lol. Thanks. No I do not, graphs and slopes and equations confuse me.

OpenStudy (mrnood):

I'm afraid I will not give you the answers A straight line graph tells you the value for Y (on the vertical axis) for each value of X on the horizontal so if y= x/4 then if x = 1 y= 1/4 if x = 2 y= 2/4 if x= 3 y = 3/4 You need to look at you text for the general form of a straight line equation y=mx+c when you have read that text then you will be able to answer the question The graph below uses the equation you have been given.. above